OpenStudy (anonymous):

I just can't solve these three problems! : ( I am isolating the x first, then substituting the x into the second part, but then I get mixed up on what to do next. Any help or hints are appreciated!! Use the substitution to solve this system of equations. 7. {y+2x=11 {x+y=5 10. {3x +y =-4 {1/2x+y =6 11. {x+2y=2 {2x+3y=-1 Thank you!

OpenStudy (anonymous):

which one were you solving?

OpenStudy (anonymous):

I was attempting to solve all three of those. :/

OpenStudy (anonymous):

hmm for 1. you have x = 5 - y <--sub that into the first equation 2. isolating x is not easy here so it's preferrable to use y. y = -4 - 3x <--sub that into the second equation 3. x = 2 - 2y <--sub that into the second equation remember to always find the easy way out. don't stick too much to the rules

OpenStudy (anonymous):

Oh! Great. Let me try that and post my answers. =)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

2y = 11/5 is what I have so far Is that okay? Then I would isolate the y by dividing 2, correct? I'm sorry, math is my worst...

OpenStudy (anonymous):

which number is this?

OpenStudy (anonymous):

The first one =)

OpenStudy (anonymous):

Sorry!! Woops I just saw my mistake!! Ahhh

OpenStudy (anonymous):

nice. i was about to point it out

OpenStudy (anonymous):

Okie dokie I got y=-1 fingers crossed! I would then use that y and input it into an original equation to get x right?

OpenStudy (anonymous):

x = 5 - y y + 2(5-y) = 11 y + 10 - 2y = 11 -y = 1 so y = -1 right

OpenStudy (anonymous):

Yay! Awesome

OpenStudy (anonymous):

yes to your second question too

OpenStudy (anonymous):

Okie dokes. Can I input y to any of the two equations?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Wow thanks!!! I should be fine on the other two problems! Thank you very much!